It takes1.5 for Tim to mow the lawn.linda can mow the same lawn in 2 hours.how long will it take john and Linda,work together to mow the lawn?
Answer by Cromlix(4381)

Tim's rate per hr = 1/1.5
Linda's rate per hr = 1/2
Together 1/1.5 + 1/2 = 3.5/3
Time to do job together
3/3.5 = 0.86 hr
or 51.4 minutes.
Hope this helps. :-)